1 4 Find the row space and null space of A= 1 0 2 2 1 -4 - 1 -2 -8
Answer #1

Find the Row space and mult space of A.- 4 2 Sol! A 4 1 1 344 Rat Rg+2 R1 1 4 0 which is in echton soven, Row Space of a, denfrom x+41W-2) + 2 2 +6 = 0 x+4 w-u2 + 2ztwo x+5w-22 = 0 x=22-5w 22-sw - 2 Z i. Null Space N(A) = { 22-5w W-z z W, ZERFind the Row space and mult space of A.- 4 2 Sol! A 4 1 1 344 Rat Rg+2 R1 1 4 0 which is in echton soven, Row Space of a, den